# Enhancing Feature Diversity Boosts Channel-Adaptive Vision Transformers

### v1

@@ 5 Conclusion
In this paper, we present DiChaViT, a model aimed at enhancing feature diversity and robustness in MCI-ViTs. First, we introduce Diverse Channel Sampling, a novel channel sampling strategy that encourages the selection of more distinct channel sets during training, thereby promoting feature diversity. Additionally, DiChaViT incorporates regularization and initialization techniques to diversify the features learned in MCI-ViTs. Our experiments demonstrate a $1.5-5.0$% point improvement over state-of-the-art methods on satellite and microscopy imaging datasets. DiChaViT represents a promising advancement in addressing the challenges associated with MCI, paving the way for more effective MCI-ViT models.
Broader Impacts and limitations. The development of DiChaViT represents an advancement in MCI, with potential positive impacts such as improved medical diagnosis and accelerated healthcare research. Additionally, its versatility in satellite imaging holds promise for environmental monitoring. However, there are also potential negative impacts, including the risk of bad actors using this research to develop harmful applications, such as invasive surveillance systems. This highlights the importance of ethical considerations and responsible deployment. One of the limitations of our work is that it is not designed to handle novel channels. Investigating techniques to adapt to new channels at test time is a promising research direction in MCI. In addition, our approach requires extra hyperparameter tuning, which may necessitate some additional compute resources.
@@
This material is based upon work supported, in part, by the NSF under award DBI-2134696. Any opinions, findings, and conclusions or recommendations are those of the author(s) and do not necessarily reflect the views of the supporting agencies.
@@ Appendix A Implementation details
CHAMMI dataset [14]. The goal of CHAMMI is to train a model to learn the feature representation for the input image. Thus, we use the [CLS] token at the final layer as the feature representation and train the model to minimize the proxy loss [52]. We then evaluate the model on various tasks following the evaluation code provided by the authors, in which a 1-Nearest Neighbour classifier is used to predict the macro-average F1-score for each task separately  [^6] https://github.com/broadinstitute/MorphEm. The channel-adaptive interfaces are adapted from the author’s implementation code  [^7] https://github.com/chaudatascience/channel_adaptive_models. Besides the model, we incorporate the same data augmentation as introduced by the authors, such as thin-plate-spline (TPS) transformations [56]. We train each model for $60$ epochs with a learning rate of $0.00004$, and a batch size of 64.
JUMP-CP [12] and So2Sat [17] datasets. Following Bao et al. [18], the learning rate is warmed up for the initial $10$ epochs, peaking at $0.0005$ after which it will gradually decay to $10^{-6}$ following a cosine scheduler. We also apply a weight decay of $0.04$ to the weight parameters, excluding the bias and normalization terms to mitigate overfitting. Additionally, we use the same data augmentation as used in the code provided by the authors. To get the final prediction, we pass the Transformer encoder’s representation for the [CLS] token into a classifier head to predict the probability of each class. We train each model for $100$ epochs, with a batch size of $64$ on JUMP-CP, and $128$ on So2Sat. We adapt the code provided by the authors [18] for the baselines in our work  [^8] https://github.com/insitro/ChannelViT.
@@ Appendix B Additional experimental results
Extended main results. Table 6 shows an extension of the main resulting table (Table 1) where we include CNN-based (ConvNeXt backbone [8]) models from [14]. To ensure a fair comparison, we adjust the number of layers in these CNN-based models so that all models in Table 6 have approximately $21$M parameters. We can observe that in general, DiChaViT outperforms CNN-based and ViT-based models on the three datasets.
@@
DiChaViT (ours) ViT 75.69 63.67 28.98 69.24 57.98 63.36 47.76
Table 7: Different choices of channel feature $f$ in Diverse Channel Sampling (DCS) (Algorithm 1). We compare the performance when using channel tokens and patch tokens (i.e., patch images after the projection layer) to compute the similarity score for sampling. Note that while channel tokens are shared across all input images, patch tokens differ for each input image. Using channel tokens yields better performance on both So2Sat and CHAMMI.
So2Sat [17] CHAMMI [14]
Patch tokens 63.00 65.57
Channel tokens 63.36 69.68
Channel feature $f$ in DCS (Algorithm 1). Table 7 compares using channel tokens and patch tokens to compute the cosine similarity in Algorithm 1. We observe that using channel tokens gains better performance on So2Sat and CHAMMI.

### v2

@@ 5 Conclusion
In this paper, we present DiChaViT, a model aimed at enhancing feature diversity and robustness in Multi-Channel Imaging (MCI) ViTs. First, we introduce Diverse Channel Sampling, a novel channel sampling strategy that encourages the selection of more distinct channel sets during training, thereby promoting feature diversity. Additionally, DiChaViT incorporates Token Diversification Loss on the patch tokens and Channel Diversification Loss for channel tokens to further diversify the features learned in MCI-ViTs. Our experiments demonstrate a $1.5-5.0$% point improvement over state-of-the-art methods on satellite and microscopy imaging datasets. Many of our enhancements are not tied to any specific architecture and can be incorporated into new architectures as they are developed. DiChaViT represents a promising advancement in addressing the challenges associated with MCI, paving the way for more effective MCI-ViT models.
Broader Impacts and limitations. The development of DiChaViT represents an advancement in MCI, with potential positive impacts such as improved medical diagnosis and accelerated healthcare research. Additionally, its versatility in satellite imaging holds promise for environmental monitoring. However, there are also potential negative impacts, including the risk of bad actors using this research to develop harmful applications, such as invasive surveillance systems. This highlights the importance of ethical considerations and responsible deployment. One of the limitations of our work is that it is not designed to handle novel channels. Generalizing to unseen channels is challenging because it requires establishing a connection between existing and new channels. This is further complicated in the presence of domain shifts, which makes finding the informative channel weights even more difficult. Thus, investigating techniques to adapt to new channels at test time is a promising research direction in MCI. In addition, our approach requires extra hyperparameter tuning, which may necessitate additional compute resources.
@@
This material is based upon work supported, in part, by the National Science Foundation under award DBI-2134696. Any opinions, findings, and conclusions or recommendations are those of the author(s) and do not necessarily reflect the views of the supporting agencies.
@@ Appendix A Implementation details
CHAMMI dataset [14]. The goal of CHAMMI is to train a model to learn the feature representation for the input image. Thus, we use the $\mathrm{[CLS]}$ token at the final layer as the feature representation and train the model to minimize the proxy loss [52]. We then evaluate the model on various tasks following the evaluation code provided by the authors, in which a $1$-Nearest Neighbour classifier is used to predict the macro-average F1-score for each task separately  [^6] https://github.com/broadinstitute/MorphEm. The channel-adaptive interfaces are adapted from the author’s implementation code  [^7] https://github.com/chaudatascience/channel_adaptive_models. Besides the model, we incorporate the same data augmentation as introduced by the authors, such as thin-plate-spline (TPS) transformations [56]. We train each model for $60$ epochs with a learning rate of $0.00004$, and a batch size of $64$.
JUMP-CP [12] and So2Sat [17] datasets. Following Bao et al. [18], the learning rate is warmed up for the initial $10$ epochs, peaking at $0.0005$ after which it will gradually decay to $10^{-6}$ following a cosine scheduler. We also apply a weight decay of $0.04$ to the weight parameters, excluding the bias and normalization terms to mitigate overfitting. Additionally, we use the same data augmentation as used in the code provided by the authors. To get the final prediction, we pass the Transformer encoder’s representation for the $\mathrm{[CLS]}$ token into a classifier head to predict the probability of each class. We train each model for $100$ epochs, with a batch size of $64$ on JUMP-CP, and $128$ on So2Sat. We adapt the code provided by the authors [18] for the baselines in our work  [^8] https://github.com/insitro/ChannelViT.
@@ Appendix B Additional experimental results
Extended main results. Tab. 7 shows an extension of the main resulting table in the main paper (Tab. 1), where we include CNN-based (ConvNeXt backbone [8]) models from [14]. To ensure a fair comparison, we adjust the number of layers in these CNN-based models so that all models in Tab. 7 have approximately $21$M parameters. We can observe that in general, DiChaViT outperforms CNN-based and ViT-based models on the three datasets.
@@
DiChaViT (ours) ViT 75.69 63.67 28.98 69.19 57.98 63.36 47.76
Extensive ablation results on DiChaViT. Tab. 8 extends Tab. 3 in the main paper to have a better understanding of the individual effects and contributions of each of the losses. We observe that adding DCS helps improve the performance (e.g., by  $4$% on CHAMMI), and robustness of the model, especially when tested on partial channels (a boost of  $35$% on So2Sat Partial). Similarly, TDL and CDL also show improvement across the three datasets. For example, TDL improves the performance by $2.5$% on CHAMMI and $1.7$% on So2Sat on full channels.
Table 8: Extensive Ablation Studies on DiChaViT. We expanded Tab. 3 in the main paper to show the performance improvements achieved with different combinations of our components, offering more insights into the roles of each component. We report mean±$\pm$±std over three runs.
CHAMMI JUMP-CP So2Sat
Exp. Model Avg Score Full Partial Full Partial
1. ChannelViT w/o HCS (ChAda-ViT) 63.88$\pm$0.34 65.03$\pm$0.98 42.15$\pm$2.33 56.98$\pm$0.46 12.38$\pm$2.03
2. + HCS (ChannelViT) 64.90$\pm$0.75 67.51$\pm$0.35 56.49$\pm$0.53 61.03$\pm$0.17 46.16$\pm$0.40
3. + DCS 67.74$\pm$0.33 67.90$\pm$0.37 56.61$\pm$0.43 62.17$\pm$0.23 47.30$\pm$0.43
4. + TDL 66.27$\pm$0.38 65.77$\pm$0.58 43.89$\pm$1.89 58.68$\pm$0.53 15.63$\pm$5.01
5. + CDL 64.24$\pm$0.54 66.75$\pm$0.57 42.74$\pm$1.74 57.70$\pm$0.11 15.08$\pm$4.00
6. + DCS + TDL 68.07$\pm$0.44 67.66$\pm$0.28 56.87$\pm$0.78 62.20$\pm$0.18 45.74$\pm$0.42
7. + TDL + CDL 65.32$\pm$0.48 66.03$\pm$0.39 42.37$\pm$1.16 59.20$\pm$0.43 17.88$\pm$3.14
8. + DCS + CDL 67.61$\pm$0.44 68.12$\pm$0.60 56.62$\pm$0.78 62.39$\pm$0.13 46.87$\pm$0.24
9. + TDL + CDL + HCS 67.46$\pm$0.39 67.50$\pm$0.90 57.10$\pm$0.96 62.05$\pm$0.09 45.08$\pm$0.60
10. + TDL + CDL + DCS (DiChaViT) 69.66$\pm$0.43 69.19$\pm$0.47 57.98$\pm$0.41 63.36$\pm$0.11 47.76$\pm$0.23
@@
Attention scores of the $\mathrm{[CLS]}$ token to the patch tokens at different layers. Fig. 7 shows an extended version of Fig. 1(b) in the main paper, where we calculate the attention scores of the $\mathrm{[CLS]}$ token to the patch tokens at layers $4$, $8$, and $12$ (the penultimate layer), and then aggregate them by channel. This indicates that ChannelViT (top) relies more heavily on specific channels (e.g., microtubules and nucleus) for making predictions, while other channels (e.g., protein and er) are less considered. In contrast, DiChaViT (bottom) displays more evenly distributed attention scores across channels, indicating that each channel contributes more significantly to the model’s predictions.
@@
Figure 7: Attention scores of HPA channels (CHAMMI) across different layers. We compute attention scores of the $\mathrm{[CLS]}$ token to the patch tokens in a layer (layers $4$, $8$, and $12$) and aggregate them by channel. ChannelViT (top) relies on certain channels (e.g., microtubules and nucleus) to make predictions and less on other channels (e.g., protein and er). In contrast, DiChaViT demonstrates more evenly distributed attention scores across channels, suggesting that each channel contributes more to the model’s predictions.
Leave-one-channel-out at test time. In Tab. 9, we provide individual channel combination results when using seven channels (of eight) of the JUMP-CP dataset for inference. This corresponds to the details in column "7" from Tab. 2 in the main paper, representing $C^{7}_{8}=8$ different channel combinations. For each combination, we report the mean and std of the models computed over three runs. Our results demonstrate that DiChaViT gets $1-2$% better performance for each combination while also providing more stable results (i.e., smaller model variance) than baseline ChannelViT.
Table 9: Detailed performances of ChannelViT and DiChaViT on JUMP-CP in the leave-one-channel-out at test time setting. We present the details of column "7" in Tab. 2 of the main paper. DiChaViT achieves $1-2$% better performance on each combination compared with ChannelViT.
Channels at inference ChannelViT DiChaViT (ours)
$\{0,1,2,3,4,5,6\}$ 67.37$\pm$0.60 69.21$\pm$0.19
$\{0,1,2,3,4,5,7\}$ 67.20$\pm$0.59 69.06$\pm$0.20
$\{0,1,2,3,4,6,7\}$ 67.28$\pm$0.53 69.12$\pm$0.16
$\{0,1,2,3,5,6,7\}$ 58.52$\pm$0.63 59.61$\pm$0.17
$\{0,1,2,4,5,6,7\}$ 37.70$\pm$0.60 38.83$\pm$0.46
$\{0,1,3,4,5,6,7\}$ 61.90$\pm$0.48 63.28$\pm$0.31
$\{0,2,3,4,5,6,7\}$ 61.21$\pm$0.41 62.72$\pm$0.28
$\{1,2,3,4,5,6,7\}$ 61.72$\pm$0.48 63.48$\pm$0.20